On October 1, 2021, the Allegheny Corporation purchased equipment for $157,000. The estimated service life of the equipment is 10 years and the estimated residual value is $3,000. The equipment is expected to produce 350,000 units during its life.

Required:
Calculate depreciation for 2021 and 2022 using each of the following methods. Partial-year depreciation is calculated based on the number of months the asset is in service

1. Straight line

2. Double-declining balance

3. Units of production (units produced in 2021, 14,000; units produced in 2022, 29,000)

4. Sum-of-the-years’-digits
Answer #1

1.

Straight Line

2021 = (157,000 - 3,000)/10 x 3/12= 3,850

2022 = (157,000 - 3,000)/10 = 15,400

2.Double Declining Balance

2021 = (157,000 - 3,000)/10 x 2 x 3/12 = 7,850

2022 = (157,000 )/10 x 2 x 9/12 = 23,550

+ (157,000- 23,550)/10 x 2 x 3/12 = 6,280

So Total Depreciation for 2022 = 23,550 + 6,280 =29,830

3.Units of Production

2021 = (157,000 - 3,000) x 14,000/350,000 =6,160

2022 = (157,000 - 3,000) x 29,000/350,000 = 12,760

4. Sum of the years Digits Method

2021 =(157,000 - 3,000)/55 x 10 x 3/12 =7,000

2022 =(157,000 - 3,000)/55 x 10 x 9/12 = 21,000

+ (157,000 - 3,000 - 28,000)/55 x 9 x 3/12 =5154.55

SO total depreciation for2022 = 21,000 + 5154.55 = 26,154.55